### v1

# DiffClean: Diffusion-based Makeup Removal for Accurate Age Estimation

Abstract
Accurate age verification can protect underage users from unauthorized access to online platforms and e-commerce sites that provide age-restricted services. However, accurate age estimation can be confounded by several factors, including facial makeup that can induce changes to alter perceived identity and age to fool both humans and machines. In this work, we propose DiffClean which erases makeup traces using a text-guided diffusion model to defend against makeup attacks. DiffClean improves age estimation (minor vs. adult accuracy by 4.8%) and face verification (TMR by 8.9% at FMR=0.01%) over competing baselines on digitally simulated and real makeup images.
1 Introduction
Motivation. Our constant need to stay digitally connected has given rise to a wide range of online services, including social media platforms, chat rooms, and dating websites. Several online services rely on third-party selfie-based age verification services for user authorization (e.g., Yoti, Ondato, TrustStamp). But there have been reports of minors bypassing age verification checks exposing them to cyberbullying and online dangers. [^1] Children bypass age-checking measures on social media sites  [^2] Underage teens bypassing age restrictions on dating apps. In such scenarios, precise age estimation is critical in protecting underage users. Age estimation is intrinsically challenging due to complex genetic, environmental, demographic, and subjective factors. The problem is compounded when an individual puts on makeup. Facial makeup significantly alters the perception of facial features that serve as visual cues for age prediction [14, 41, 9]. The perceived age in younger people with makeup is generally overestimated, while, in older people, the predicted age is underestimated [52, 54]. Makeup can also fool automated age estimation frameworks [6]. Therefore, we need a reliable makeup-invariant age estimator.
Intuitively, we can improve the robustness of the age estimation algorithms to the presence of makeup by using “more data”. However, this immediately raises privacy concerns: we will need data on minors wearing makeup. The next strategy might be to generate synthetic makeup images for training age classifiers. But makeup can vary in terms of intensity, texture, color, region of application, and subjective preferences. So, synthetic data may not capture the full diversity of facial makeup. Thus, we propose to design a makeup removal framework to erase makeup traces digitally, and help restore the actual age. Additionally, our method also recovers identity cues confounded by makeup.
Our approach. Makeup removal can simultaneously benefit age estimation and face-based identity verification. To achieve this objective, we formulate makeup removal as an image translation task. Naturally, we adopt generative models that are adept at facial editing; specifically, seamless and controllable digital makeup transfer [30, 7, 24, 34, 56, 43, 48]. Our goal differs from previous work that focused on attribute editing and adversarial makeup transfer. We need to restore the age cues of original image that may have been perturbed after applying makeup. So, we propose DiffClean that uses a text-guided diffusion model to transform a face image with makeup to a face image without makeup using a combination of CLIP loss (for makeup detection), perceptual loss (for fidelity retention), biometric loss (for identity restoration), and age loss (for age restoration); see Fig. 1. DiffClean is efficient and can be plugged in as a pre-processing module to erase makeup traces, if present, for downstream facial analytics.
Contributions. Our main contributions are as follows:
  • We propose Diffclean, a novel text-guided diffusion-based makeup removal framework that digitally erases the traces of makeup present in face images, and improves age estimation (minor vs. adult accuracy by 4.8%) and face verification (TMR by 8.9% at FMR=0.01%) over state-of-the-art baselines.
  • We utilize text-guided CLIP loss for detecting makeup, perceptual loss for minimizing visual artifacts, and a combination of age and identity losses for restoring the facial cues affected by makeup. Our method shows improved generalizability over existing methods in terms of different makeup styles and demographic factors.
  • We demonstrate that our method works on both digitally simulated and real makeup images, thus capable of mitigating differential effects of makeup that may help protect underage users from cyber harms. We will release the codes for reproducible research.
Responsible Usage. We recognize that our application focus is a sensitive topic, particularly since any modeling or algorithmic errors may adversely impact a vulnerable demographic (minors/teenagers). We strongly advocate for the ethical use of DiffClean only to assist with facial analytics (not for the purposes of malicious image editing).
Figure 1: Overview of our method, DIFFCLEAN, that removes makeup on an example FFHQ image using text-guided diffusion model with a combination of CLIP loss, identity (ID) loss, Age loss and Perceptual (Visual) losses. Successful makeup removal results in shifting of softmax probabilities of $[p(withmakeup),p(withoutmakeup)]$ from $[0.76,0.24]$ to $[0.13,0.87]$, and reduces overestimation from 14.6 yrs to 10.4 yrs.
2 Related Work
Facial Age Prediction. Classical approaches extract handcrafted features to design classification and regression models. [28] used statistical face models generated by applying Principal Component Analysis (PCA) on a set of face images for age estimation. Texture features such as Local Binary Pattern (LBP) and its variants have been used for reliable age estimation [17, 44, 59]. Biologically-inspired features have been used in [18, 20]. Deep-learning -based approaches rely on convolutional neural networks (CNNs) and transformers for improved robustness. A multi-stream CNN was used to learn high-dimensional structured information from patches in facial images for reliable age estimation in [58]. Deep CNN was used for age estimation in unconstrained settings in [38, 53, 29]. In [36], each age was treated as a rank to design an end-to-end learning approach to apply ordinal regression problem to age estimation using deep CNN. A multi-task learning approach for heterogeneous attribute estimation was adopted in [21]. Deep Expectation (DEX) [39] formulated age estimation by converting it into a classification-regression problem. SSRNet [57] adopted a compact soft stagewise regression model that employed a coarse-to-fine strategy to perform multi-class classification. Auxiliary facial cues were used to boost reliable age estimation in [33, 60]. Deep random forests were used for age estimation in [45]. FP-age [32] used face parsing attention module to incorporate facial semantics into the age estimation. Swin Transformer with attention-based convolution (ABC) for facial age estimation was proposed in [46]. In [27], a multi-input transformer-based model, MiVOLO, was designed for age and gender prediction using facial and full body information.
Figure 2: Makeup diversity due to variations in intensity (light vs. heavy), region of application (eye vs. lip), color (pink vs. red), texture (matte vs. gloss), etc. Example images are generated by synthetic retouching the original image from CelebA-HQ[25] using Google Gemini Flash 2.0.
Age Estimation in Minors. While there is extensive research on age estimation for adults, limited work have focused on improving the age estimation performance for minor age groups. In [3], a finetuned VGG-based ‘children’ network was used for apparent age estimation of children of age from 0 to 12 years. [2] trained ResNet50 on underage facial dataset ‘VisAGe’ to improve age estimation for minors. The need for fast, accurate and robust age estimation and verification has gained traction recently. NIST Face Analysis Technology Evaluation (FATE) Age Estimation & Verification reports that subjects aged between 8 and 12 years tend to be overestimated to be 13 to 16 years old with false positives as high as $0.76$  [^3] NIST FATE Report; higher errors are reported for females than for males, specifically, in 17-year-olds [22]. Performance of facial age estimation algorithms can be affected by several factors like gender, ethnicity, image quality, facial expression, and makeup [1, 8].
Impact of Cosmetic Makeup on Facial Analytics. [51] demonstrated the impact of facial makeup on automatic face recognition. [14] conducted a comprehensive study suggesting that makeup applied to different facial regions also has different effects on age prediction and further quantified the makeup effect as Young Index (YI). [13] observed that Identification Rate (IDR) decreases with increasing amount of makeup. Taking into account the adverse effects of makeup in face verification task, [5, 19, 23] focused on improving robustness of face recognition to makeup. Authors in [6] observed that facial cosmetics (both real makeup and synthetically retouched) affect automated age estimation causing a mean absolute error of up to 5.84 years. They further indicated that lip makeup had least impact on automated age estimation, followed by eye makeup, followed by full face makeup (eye + lip + foundation). Makeup has shown to produce differential effects in apparent (manual) age estimation. 20-year-old women appeared older when wearing full face makeup while 40- and 50-year-old women appeared younger when wearing makeup [10, 8, 41, 9]. Limited work discuss the influence of makeup on facial age estimation in minors [1, 2].
Facial Makeup Transfer and Removal. Approaches for makeup transfer relied on GAN-based style-transfer networks to map reference makeup styles to target image [30, 7, 24, 34, 56, 50]. Later, diffusion models (and recently, multimodal LLMs) bridged this gap [62, 49, 40]; see examples of diverse makeup editing in Fig. 2. Makeup removal, on the other hand, poses a more challenging task where the goal is to reveal the face under makeup without adversely affecting the biometric and perceptual features. To assist with face verification, [31] integrated two adversarial networks, both at pixel and feature levels to generate identity-preserving synthetic non-makeup faces, while, [61] replaced convolution layers in GAN with resnet blocks to design makeup removal network. [40] proposed an one-for-all diffusion-based framework for a wide array of makeup tasks including makeup transfer and removal by modeling each as a domain translation problem. CLIP2Protect [43] and DiffAM [48] use adversarial makeup transfer for privacy with capability to generate ‘no makeup’ faces.
3 Proposed Method
Makeup has an adverse effect on facial analytics, specifically, exposing minors to online harms by breaching automated age-based regulation checks. Therefore, addressing this issue in a principled way will have a positive societal impact. One workaround is to collect representative data of minors wearing makeup for improving robustness; this poses strong privacy concerns. So, we propose a generative model, Diffclean, that is designed to mitigate the age variations arising due to makeup. There are two challenges to designing this network: (1) we need a robust makeup style generator to synthesize data needed for training thus circumventing data collection; (2) we need a robust age estimator to serve as a proxy auxiliary classifier to guide DiffClean to restore age-specific cues.
Figure 3: Synthetic makeup transfer results produced using EleGANt [56] on example images from UTKFace. The first column indicates the original image, the second column indicates the makeup style reference, and the third column indicates the images after makeup transfer.
For the makeup style generator, we used an existing network, EleGANt [56] that can perform locally-editable makeup transfer; see Fig. 3. Age estimation algorithms typically perform poorly on lower age groups which are most vulnerable to overestimation due to the application of makeup. Therefore, we first need an auxiliary age estimator that has been specifically fine-tuned on our target age groups (10-19 years). For the proxy age classifier, we used SSRNet [57] fine-tuned on UTKFace [63] dataset with weighted self-adjusting smoothed-$\mathcal{L}_{1}$ loss to aggressively penalize misclassifications in target age groups.
DiffClean uses the ablated diffusion model (ADM) with guided diffusion from [12] with improved U-Net architecture containing increased attention heads at multiple resolutions, BigGAN residual stacks for upsampling and downsampling, deeper networks and rescaled residual connections; refer to  [12] for details. We used text prompts face with makeup and face without makeup for text-guided CLIP directional loss following [48]. The CLIP loss helps remove makeup traces conditioned on identity, age and perceptual constraints. For training the diffusion model, we sampled images from the UTKFace [63] dataset that consists of original face images and applied EleGANt to generate (w/o makeup, w/ makeup) pairs. We evaluated our trained DiffClean on FFHQ [26] dataset and used the exact age values provided by [11] in conjunction with fine-tuned SSRNet to ensure that the predicted age of the image after makeup removal is similar to the input image with makeup.
Methodology. We describe the terminology and notation here. Original/Clean image (without makeup): $\textbf{I}_{o}$, Makeup image (with real or synthetic makeup), $\textbf{I}_{m}$, and Makeup removed image, $\textbf{I}_{g}$. Consider $\mathcal{M}$ as a real or synthetic makeup application operator, while $\mathcal{M}^{-1}$ represents the makeup removal operator. Thus, $\textbf{I}_{o}\xrightarrow[]{\mathcal{M}}\textbf{I}_{m}$, while $\textbf{I}_{m}\xrightarrow[]{\mathcal{M}^{-1}}\textbf{I}_{g}$. The goal is to model $\mathcal{M}^{-1}$ so that we can recover the clean image, $\textbf{I}_{g}\approx\textbf{I}_{o}$. Our proposed framework DiffClean approximates the role of $\mathcal{M}^{-1}$.
Architecture: The makeup removal framework ($\mathcal{M}^{-1}$) used in this work is a generative model which is a pre-trained ADM [12] that adds noise to the input image (with makeup), $\textbf{I}_{m}$, to generate the noisy latent image during forward diffusion. The deterministic reverse diffusion through DDIM produces the makeup removed output, $\textbf{I}_{g}$ supervised by CLIP, biometric identity, age and perceptual losses.
Loss functions:
Proxy Age Estimator: The age estimator, SSRNet is fine-tuned using a self-adjusting smoothed-$\mathcal{L}_{1}$ loss introduced in [15] and adaptively weighted for different age groups as follows:
$\mathcal{L}_{WSL}=\begin{cases}3.0*\mathcal{L}_{SL},\text{if age $\{10-29\}$ yrs.}\\ 1.0*\mathcal{L}_{SL},\text{otherwise.}\end{cases}$ (1)
where,
$\mathcal{L}_{SL}(a,\hat{a})=\begin{cases}0.5\frac{(a-\hat{a}).^{2}}{\beta},&\text{if $a-\hat{a}<\beta$}\\ |a-\hat{a}|-0.5*\beta,&\text{otherwise}\end{cases}$
Here, $a$ is the ground truth age, $\hat{a}$ is the predicted age, and the parameter $\beta$ is adapted based on the running mean and variance. Note we ablated with different losses such as $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and Huber loss functions but empirically observed weighted self-adjusting smooth $\mathcal{L}_{1}$ loss to perform the best.
DiffClean: During training, we used four losses to fine-tune the diffusion model as follows.
CLIP-based Makeup Loss: The transition from makeup domain to non-makeup domain is guided via textual prompts ‘face without makeup’ and ‘face with makeup’ in CLIP space following [48]. The loss is one minus the cosine similarity between the image and text embeddings as follows:
$\mathcal{L}_{clip}=1-\frac{\Delta\textbf{I}.\Delta\textbf{T}}{\left\lVert\Delta\textbf{I}\right\rVert\left\lVert\Delta\textbf{T}\right\rVert}$ (2)
where $\Delta\textbf{I}=\mathcal{E}_{\textbf{I}}(\textbf{I}_{g})-\mathcal{E}_{\textbf{I}}(\textbf{I}_{m})$ and $\Delta\textbf{T}=\mathcal{E}_{\textbf{T}}(t_{\text{face w/o makeup}})-\mathcal{E}_{\textbf{T}}(t_{\text{face w/ makeup}})$. Here, $t_{\text{face w/o makeup}}$ is textual prompt after makeup removal and $t_{\text{face w/ makeup}}$ is textual prompt before makeup removal. Here, $\mathcal{E}_{\textbf{I}}$ and $\mathcal{E}_{\textbf{T}}$ are CLIP image and text encoders, respectively.
Biometric identity Loss: Our goal involves identity restoration of the makeup removed image with respect to the original image (without makeup). During training with synthetic makeup pairs, we have access to ($\textbf{I}_{o}$, $\textbf{I}_{m}$) to generate $\textbf{I}_{g}$. So, we compute weighted pairwise biometric identity losses as $\mathcal{L}_{og}$ and $\mathcal{L}_{mg}$. This is different than [48] that uses a single identity loss between input and output. The final identity loss is as follows:
$\mathcal{L}_{id}=\lambda_{1}\mathcal{L}_{og}+\lambda_{2}\mathcal{L}_{mg}$ (3)
where $\mathcal{L}_{\cdot,\cdot}$ represents the cosine distance between the extracted features using the ResNet ArcFace from the InsightFace library, with weights $\lambda_{1}=0.75$ and $\lambda_{2}=0.25$.
Perceptual Loss: To preserve the perceptual features, we use the perceptual loss $\mathcal{L}_{pips}$ with VGG [47] network and $\mathcal{L}_{1}$ loss between generated makeup-removed image $\textbf{I}_{g}$ and input image with makeup $\textbf{I}_{m}$.
Age Loss: Restoration of age cues requires explicit fine-tuning of the makeup removal module with age loss. The age loss is computed as the smoothed-$\mathcal{L}_{1}$ loss between the the ground truth age, $a_{i}$ of the original image, $\textbf{I}_{o}$, and the predicted age, $\hat{a_{i}}$ of the makeup removed image, $\textbf{I}_{g}$. We used the fine-tuned SSRNet model to predict $\hat{a_{i}}$. Note age loss needs a specific value, whereas, FFHQ-Aging provides age groups. So, we used the user metadata (age labels) provided by [11]. We introduce the SSRNet age loss as follows:
$\mathcal{L}_{\text{SSRNet-age}}=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{SL}(a_{i},\hat{a}_{i})$ (4)
[…]
$\mathcal{L}_{\text{CLIP-age}}=1-\frac{\textbf{I}_{g}.\textbf{t}_{age}}{\left\lVert\textbf{I}_{g}\right\rVert\left\lVert\textbf{t}_{age}\right\rVert}$ (5)
Total loss for makeup removal: Combining all the above losses, the final loss for DiffClean is as follows:
$\mathcal{L}_{total}=\lambda_{clip}\mathcal{L}_{clip}+\lambda_{id}\mathcal{L}_{id}+\lambda_{lpips}\mathcal{L}_{lpips}+\lambda_{L1}\mathcal{L}_{L1}+\lambda_{age}\mathcal{L}_{age}$ (6)
where loss weights are tuned as $\lambda_{clip}=5.0$, $\lambda_{id}=1.0$, $\lambda_{lpips}=5.0$, $\lambda_{L1}=2.0$, and $\lambda_{age}=0.5$ for SSRNet age loss and $\lambda_{age}=5.0$ for CLIP age loss. Thus, we design two models: DiffClean with SSRNet age loss ($\mathcal{L}_{age}=\mathcal{L}_{\text{SSRNet-age}}$) and DiffClean with CLIP age loss ($\mathcal{L}_{age}=\mathcal{L}_{\text{CLIP-age}}$).
During testing, our model can accept any face image (w/o makeup or w/ real/synthetic makeup) as input. We perform age estimation using MiVOLO [27] (while training uses SSRNet) and MobileFace and FaceNet for identity verification (while training uses ArcFace IRSE50).
4 Experiments and Analysis
4.1 Datasets
Training/Fine-tuning: We used the UTKFace [63] dataset to fine-tune SSRNet for age estimation with a training set of 15,364 images and a testing set of 6,701 images. The original set of 20K images was split into training and testing splits (70:30) after filtering out images with ages $\geq$70 yrs. We adopted a two-step approach for fine-tuning DiffClean for makeup removal. First, we used 300 images (200 training, 100 validation) from MT dataset [30] following the guidelines in [48]. We further refined the model by fine-tuning it additionally on 300 images (200 training, 100 validation) from the UTKFace dataset after digitally applying makeup using EleGANt [56] with the combined loss. Thus, our training set, $D_{tr}$ consisted of 600 images in total. We used five reference images [30] to simulate makeup styles while using EleGANt. Our focus is makeup removal which is a many-to-one task, i.e., translate from any makeup style to recover the clean non-makeup image. So, we used a few representative synthetic makeup styles and tested the robustness of our solution in real-world settings.
Evaluation: We tested our proposed DiffClean on both synthetic makeup and real-world makeup. For synthetic makeup, we sampled 2,556 images from the Flickr-Faces-HQ Dataset (FFHQ) [26] by sampling 284 images from each of the 9 age groups (excluding $\geq 70$ yrs.) using the age bins provided by FFHQ-Aging [37]. Next, we applied EleGANt to curate synthetic makeup dataset to assess the performance of our makeup removal framework. For real-world makeup, we used 3,000 images from the BeautyFace [55] dataset and 355 images from the LADN dataset [16] for testing generalizability of our method.
Figure 4: Makeup removal results generated by different methods. (First column:) Original images. (Second column:) Baseline outputs from CLIP2Protect [43] and DiffAM [48]. (Third column:) Our outputs from DiffClean with SSRNet and CLIP age loss. Our method significantly removes makeup (DiffAM retains makeup) without visual artifacts (CLIP2Protect changes hair color and facial features).
4.2 Metrics
We evaluated the model for age restoration and identity retention. For age estimation, we report the Mean Absolute Error (MAE) between ground truth and predicted age values (lower is better); Age group accuracy to assess if the predicted age is correctly grouped into age bins (higher is better); Minor/Adult Accuracy to evaluate if the predicted age $\geq 18$ to classify as adult, otherwise, minor (higher is better). For identity verification, we report the True Match Rate (TMR) at a False Match Rate (FMR) of 0.01% and further show the ROC curves.
4.3 Implementation Details
The SSRNet model was fine-tuned with batch size 50 with Adam optimizer with weight decay as 1e-4, cosine annealing scheduler and learning rate set to 1e-3 for 200 epochs with early stopping. The input image size is 64$\times$64. The makeup removal module takes the makeup image, $\textbf{I}_{m}$, of size 256$\times$256 as input, and uses 40 DDIM inversion steps with 6 DDIM sampling steps and (total time steps=80), and noise scale for latent image generation is controlled by cosine beta scheduler[35]. We used the Adam optimizer with a learning rate of 4e-3. The model is fine-tuned for 5 epochs. All experiments were carried out on a single NVIDIA A100 GPU. Hyperparameters in Eqn. 6 were obtained using grid search on the validation set.
4.4 Baselines
We compare DiffClean with 2 baselines: (1) CLIP2Protect [43], and (2) DiffAM [48]. To implement CLIP2Protect as a baseline, we used the available pretrained checkpoint with ‘no makeup’ text prompt and set the target identity to the source identity. For adversarial makeup transfer, we used single target identity and ‘Matte’ makeup style as it achieved highest success [43]. To implement DiffAM as a baseline (reproducing paper results), we needed to fine-tune the pretrained model provided by the authors [30] on MT dataset. For adversarial makeup transfer experiments, we used the checkpoint provided by the authors [48], and the default parameters corresponding to the target model, reference image, and target identity.
5 Results and Discussion
We present qualitative results for subjective evaluation and quantitative results for objective evaluation. We compare with the baselines in terms of age estimation and identity retention. See Fig. 4 for visualizing outputs of images before and after makeup removal and the corresponding outputs produced by the baselines.
5.1 Age estimation
We report the results of the predicted age estimated by MiVOLO after makeup removal in Table 1. We compare the predicted age of ‘Makeup’ images simulated using EleGANt and compare with the age labels of the original FFHQ images (without makeup) and compute the minor vs adult accuracy (%), age group accuracy (%), and mean absolute error (MAE). In Table 1, our method DiffClean outperforms both baselines, CLIP2Protect and DiffAM, and can successfully restore the age cues disrupted due to makeup by lowering the MAE (to 5.71) and improving the overall age group prediction accuracy (to 37%). Our objective focused on minor age groups, and results indicate that we achieve superior minor vs adult age classification of 88.6%.
Figure 5: Examples of makeup removal on sample images from CelebA-HQ [25] and CACD [4] datasets while mitigating differential effects of makeup. Each column consists of pairs of images before makeup removal (left) and after makeup removal (right). In the first column, DiffClean results in lower predicted ages, while in the second column, DiffClean results in similar predicted ages, and finally, in the third column, DiffClean results in higher predicted ages than the original.
Makeup exhibits a differential effect on facial age. It may result in a decrease or increase in apparent age, and in some cases, may not affect perceived age. Our method does not arbitrarily reduce predicted age in all cases; instead, it strategically erases the makeup traces to restore the original age-specific cues. Fig. 5 clearly depicts the differential age restoration where the predicted age after makeup removal may be lower, equal, or higher than the estimated age on makeup images.
Table 1: Results of age estimation in terms of minor vs adult accuracy (%) $\uparrow$, age group accuracy (%) $\uparrow$ and mean abolsute error (MAE) $\downarrow$. Our method DiffClean outperforms baselines and improves overall age estimation with respect to makeup images.
Metrics Makeup images
CLIP2Protect
(‘no makeup’) prompt
DiffAM
DiffClean (Ours)
SSRNet age loss
DiffClean (Ours)
with CLIP age loss
Minor/Adult Acc (%) 82.8 83.8 85.5 87.8 88.6
Age group Acc (%) 34.4 33.3 34.5 36.8 37.0
MAE 6.56 6.54 6.32 5.76 5.71
FaceNet
MobileFace
Methods FR model
FaceNet MobileFace
Original 74.5 79.1
CLIP2Protect 67.8 ($\downarrow$) 75.3 ($\downarrow$)
DiffAM 69.1 ($\downarrow$) 78.2 ($\downarrow$)
DiffClean (Ours-SSRNet) 75.8 ($\uparrow$) 84.2 ($\uparrow$)
DiffClean (Ours-CLIP) 75.5 ($\uparrow$) 82.4 ($\uparrow$)
Figure 6: (Top): ROC curve with FaceNet. (Middle): ROC curve with MobileFace. (Bottom): Biometric matching in terms of TMR (%) @FMR = 0.01% (higher is better) with FaceNet and MobileFace matchers. Our DiffClean improves overall performance.
5.2 Identity verification
We present the results of biometric identity verification in Fig. 6 using FaceNet and MobileFace face recognition (FR) models at FMR=0.01%. We compute the ‘Original’ performance by comparing the original FFHQ face images (without makeup) with images simulated using EleGANt (with makeup). Baseline performances are computed by comparing the original FFHQ face images with outputs of respective baselines (makeup removed images). Finally, we report the biometric matching of our proposed method by comparing the original FFHQ face images with outputs of DiffClean (makeup removed images). We observe that makeup removal improves face verification as initially hypothesized. The True Match Rate (TMR) improves by 1.3% with FaceNet and 5.1% with MobileFace compared to ‘Original’ performance, while TMR improves by 8.0% with FaceNet and 8.9% with MobileFace compared to the baselines. Note both baselines degrade identity cues.
5.3 Real-world makeup generalizability
Real-world examples only include image with makeup, but not the actual age label or the original image before makeup. So, it is necessary to evaluate the underestimation and overestimation in predicted ages after makeup removal. We used 3,000 makeup images from the BeautyFace [55] dataset and 355 images from the LADN [16] dataset. We tested DiffClean (with SSRNet and CLIP-based age loss) on these images, and used MiVOLO to predict the age before and after makeup removal; see Fig. 7. We observe that DiffClean lowers both overestimation and underestimation errors induced by makeup. Results indicate that our method reasonably generalizes across a broad spectrum of real-world makeup styles. Refer to Tables 2 and 3 for detailed statistics. We further report the confidence interval (CI) and margin of error on the BeautyFace dataset using the student’s t-distribution at a 95% confidence level. CI = [2.88, 3.09] with margin of error = 0.10 on the underestimation errors, and CI = [2.75, 2.89] with margin of error of 0.07 on the overestimation errors. This shows that our method can reduce the effects of makeup on estimated age by ±$\pm 3$ years on average.
Table 2: Statistics of predicted ages on the real-world images from the BeautyFace dataset before and after makeup removal.
Before makeup
removal
After makeup
removal
$\mu$$\pm$ $\sigma$ $\mu$$\pm$ $\sigma$
Under
estimation
$\mu$$\pm$ $\sigma$
Over
estimation
$\mu$$\pm$ $\sigma$
22.38 $\pm$ 3.36
DiffClean
21.28 $\pm$ 4.58 2.99 $\pm$ 2.92 2.79 $\pm$ 1.96
DiffClean
21.34 $\pm$ 4.66 3.14 $\pm$ 2.97 2.82 $\pm$ 1.95
Table 3: Statistics of predicted ages on the real-world images from the LADN dataset before and after makeup removal.
Before makeup
removal
After makeup
removal
$\mu$$\pm$ $\sigma$ $\mu$$\pm$ $\sigma$
Under
estimation
$\mu$$\pm$ $\sigma$
Over
estimation
$\mu$$\pm$ $\sigma$
24.64 $\pm$ 3.88
DiffClean
22.59 $\pm$ 4.47 1.60 $\pm$ 1.11 3.02 $\pm$ 1.90
DiffClean
22.55 $\pm$ 4.51 1.62 $\pm$ 1.11 3.11 $\pm$ 1.97
Figure 7: Results of DiffClean on real-world makeup images from LADN [16] (left) and BeautyFace [55] (right) datasets. Our method is capable of reducing makeup induced overestimation errors in predicted age (indicated in white) in practical settings.
5.4 Ablation study
Training Dataset Diversity. We fine-tuned on 600 images (300 from MT dataset with smoky-eyes, flashy, retro, Korean and Japanese makeup styles + 300 from UTKFace with EleGANt-based makeup transfer with five reference styles and adaptive interpolated shading). To further assess training data diversity, we augmented our training set, $D_{tr}$, and report MAE on validation set, as follows: 4.42 ($\lvert D_{tr}\rvert=600$), 4.55 ($\lvert D_{tr}\rvert=1,200$) and 4.68 ($\lvert D_{tr}\rvert=1,600$). We observe that MAE is relatively stable across different training dataset scale.
Influence of each loss component. We report MAE ($\downarrow$) on the validation set by fine-tuning separate models with loss components added incrementally, as follows: 9.28 ($\mathcal{L}_{clip}$), 7.43 ( $\mathcal{L}_{clip}+\mathcal{L}_{id}$), 4.70 ($\mathcal{L}_{clip}+\mathcal{L}_{id}+\mathcal{L}_{pips}$), and 4.42 ($\mathcal{L}_{clip}+\mathcal{L}_{id}+\mathcal{L}_{pips}+\mathcal{L}_{SSRNet-age}$). Evidently, the combined losses improve the overall performance.
Demographic fairness analysis. We conducted demographic fairness analysis on FFHQ using DeepFace library [42]; see Table LABEL:Tab:fair. Note FFHQ has more White images than Black and Indian images, but a balanced gender distribution. We observe that DiffClean does not exhibit demographic disparity.
Inference time. We compared the inference time of makeup removal using our method with CLIP2Protect. We observed that our method is 15$\times$ faster; Ours: 2 secs. and CLIP2Protect: 30 secs.
Table 4: Demographic fairness analysis on FFHQ using our DiffClean SSRNet/CLIP-age loss models in terms of MAE. Values in parentheses indicate count after being processed by DeepFace.
Race
White
(1,420)
Black
Asian
Hispanic
Indian
Middle
Eastern (129)
MAE 5.9/5.9 4.2/4.3 5.3/5.3 5.5/5.4 6.9/6.8 5.3/5.2
Gender Male (1,188) Female (1,168)
MAE 5.6/5.5 5.9/5.8
6 Conclusion
In this work, we investigate the issue of inadequacies of age estimators in scenarios where a person wears makeup. This poses a serious concern when the individual is underage or minor and can bypass age verification checks on adult-themed online services. To address these limitations, we designed DiffClean, a diffusion-based makeup removal framework that can easily remove makeup traces, and can be deployed as a ready-to-plugin pre-processing module for downstream facial analytics. Using the guidance from CLIP space, we optimized losses for identity preservation and perceptual fidelity retention to generate makeup-removed face images while restoring age-relevant features via fine-tuned auxiliary age predictor. Evaluations show effectiveness of our approach over competing baselines in terms of age estimation and identity verification while exhibiting robust generalizability across real-world makeup styles.